# Can a MISL Fly? Analysis and Ingredients for Mutual Information Skill Learning

### v2

@@ 1 Introduction
Self-supervised learning has had a large impact on areas of machine learning ranging from audio processing (Oord et al., 2016; 2018) or computer vision (Radford et al., 2021; Chen et al., 2020) to natural language processing (Devlin et al., 2019; Radford & Narasimhan, 2018; Radford et al., 2019; Brown, 2020). In the reinforcement learning (RL) domain, the “right” recipe to apply self-supervised learning is not yet clear. Several self-supervised methods for RL directly apply off-the-shelf methods from other domains such as masked autoencoding (Liu et al., 2022), but have achieved limited success so far. Other methods design self-supervised routines more specifically built for the RL setting (Burda et al., 2019; Pathak et al., 2017; Eysenbach et al., 2019; Sharma et al., 2020; Pong et al., 2020). We will focus on the skill learning methods, which aim to learn a set of diverse and distinguishable behaviors (skills) without an external reward function. This objective is typically formulated as maximizing the mutual information between skills and states (Gregor et al., 2016; Eysenbach et al., 2019), namely mutual information skill learning (MISL). However, some promising recent advances in skill learning methods build on other intuitions such as Lipschitz constraints (Park et al., 2022) or transition distances (Park et al., 2023). This paper focuses on determining whether the good performance of those recent methods can still be explained within the well-studied framework of mutual information maximization.
@@
Building upon our new interpretations of METRA, we propose a simpler and competitive MISL algorithm called Contrastive Successor Features (CSF). First, CSF learns state representations by directly optimizing a contrastive lower bound on mutual information, preventing the dual gradient descent procedure adopted by METRA. Second, while any off-the-shelf RL algorithm (e.g. SAC (Haarnoja et al., 2018)) is applicable, CSF instead learns a policy by leveraging successor features of linear rewards defined by the learned representations. Experiments on six continuous control tasks show that CSF is comparable with METRA, as evaluated on exploration performance and on downstream tasks. Furthermore, ablation studies suggest that rewards derived from the information bottleneck as well as a specific parameterization of representations are key for good performance.
@@
Figure 1: From METRA to MISL. (Left) METRA argues optimizing a Wasserstein distance is superior to using mutual information. (Right) Through careful analysis, we show METRA still bears striking similarities to MISL algorithms, which allows us to develop a new MISL algorithm (CSF) that matches the performance of METRA while retaining the theoretical properties associated with MI maximization.
@@ 2 Related Work
Unsupervised skill discovery. Our work builds upon prior methods that perform unsupervised skill discovery. Prior work has achieved this aim by maximizing lower bounds (Tschannen et al., 2020; Poole et al., 2019) of different mutual information formulations, including diverse and distinguishable skill-conditioned trajectories (Li et al., 2023; Eysenbach et al., 2019; Hansen et al., 2020; Laskin et al., 2022; Strouse et al., 2022), intrinsic empowerment (Mohamed & Jimenez Rezende, 2015; Choi et al., 2021), distinguishable termination states (Gregor et al., 2016; Warde-Farley et al., 2019; Baumli et al., 2021), entropy bonus (Florensa et al., 2016; Lee et al., 2019; Shafiullah & Pinto, 2022), predictable transitions (Sharma et al., 2020; Campos et al., 2020), etc. Among those prior methods, perhaps the most related works are CIC (Laskin et al., 2022) and VISR (Hansen et al., 2020). We will discuss the difference between them and our method in Sec. 5. Another line of unsupervised skill learning methods utilize ideas other than mutual information maximization, such as Lipschitz constraints (Park et al., 2022), MDP abstraction (Park & Levine, 2023), model learning (Park et al., 2023), and Wasserstein distance (Park et al., 2024; He et al., 2022). Our work will analyze the state-of-the-art method named METRA (Park et al., 2024) that builds on the Wasserstein dependency measure (Ozair et al., 2019), provide an alternative explanation under the well-studied MISL framework, and ultimately develop a simpler method.
@@
Contrastive learning. Contrastive learning has achieved great success for representation learning in natural language processing and computer vision (Radford et al., 2021; Chen et al., 2020; Gao et al., 2021; Sohn, 2016; Chopra et al., 2005; Oord et al., 2018; Gutmann & Hyvärinen, 2010; Ma & Collins, 2018; Tschannen et al., 2020). These methods aim to push together the representations of positive pairs drawn from the joint distribution, while pushing away the representations of negative pairs drawn from the marginals (Ma & Collins, 2018; Oord et al., 2018). In the domain of RL, contrastive learning has been used to define auxiliary representation learning objective for control (Laskin et al., 2020; Yarats et al., 2021), solve goal-conditioned RL problems (Zheng et al., 2024; Eysenbach et al., 2022; 2020; Ma et al., 2023), and derive representations for skill discovery (Laskin et al., 2022). Prior work has also provided theoretical analysis for these methods from the perspective of mutual information maximization (Poole et al., 2019; Tschannen et al., 2020) and the geometry of learned representations (Wang & Isola, 2020). Our work will combine insights from both angles to analyze METRA and show its relationship to contrastive learning, resulting in a new skill learning method.
@@
Successor features. Our work builds on successor representations (Dayan, 1993), which encode the discounted state occupancy measure of policies. Prior work has shown these representations can be learned on high-dimensional tasks (Kulkarni et al., 2016; Zhang et al., 2017) and help with transfer learning (Barreto et al., 2017). When combined with universal value function approximators (Schaul et al., 2015), these representations generalize to universal successor features, which estimates a value function for any reward under any policy (Borsa et al., 2019). While prior methods have combined successor feature learning with mutual information skill discovery for fast task inference (Hansen et al., 2020; Liu & Abbeel, 2021), we instead use successor features to replace $Q$ estimation after learning state representations (Sec. 5).
@@ 3 Preliminaries
where $p^{\pi}(s_{+}=s\mid z)$ is the discounted state occupancy measure (Ho & Ermon, 2016; Nachum et al., 2019; Eysenbach et al., 2022; Zheng et al., 2024) of policy $\pi$ conditioned on skill $z$, and $p^{\pi}(s^{\prime}\mid s,z)$ is the state transition probability given policy $\pi$ and skill $z$. This optimization problem can be casted into an iterative min-max optimization problem by first choosing a variational distribution $q(z\mid s,s^{\prime})$ to fit the historical posterior $p^{\beta}(z\mid s,s^{\prime})$, which is an approximation of $p^{\pi}(z\mid s,s^{\prime})$, and then choosing policy $\pi$ to maximize discounted return defined by the intrinsic reward $\log q(z\mid s,s^{\prime})$:
@@
where $k$ indicates the number of updates. See Appendix A.1 for detailed discussion.
@@
METRA. Maximizing the mutual information between states and latent skills $I(S;Z)$ only encourages an agent to find discriminable skills, while the algorithm might fail to prioritize state space coverage (Park et al., 2024; 2022). A prior state-of-the-art method METRA (Park et al., 2024) proposes to solve this problem by learning representations of states $\phi:{\mathcal{S}}\mapsto\mathbb{R}^{d}$ via maximizing the Wasserstein dependency measure (WDM) (Ozair et al., 2019) between states and skills $I_{\mathcal{W}}(S;Z)$. Specifically, METRA chooses to enforce the 1-Lipschitz continuity of $\phi$ under the temporal distance metric, resulting in a constrained optimization problem for $\phi$:
@@ 4 Understanding the Prior Method
In this section we reinterpret METRA through the lens of MISL, showing that:
@@
    The METRA actor objective is equivalent to a mutual information lower bound plus an extra term. This extra term is related to an information bottleneck (Tishby et al., 2000; Alemi et al., 2017) and our experiments will show it is important for exploration. See Sec. 4.2.
@@ 4.1 Connecting METRA’s Representation Objective and Contrastive Learning
Our understanding of METRA starts by interpreting the representation objective of METRA as a contrastive loss. This interpretation proceeds by two steps. First, we focus on understanding the actual representation objective of METRA, aiming to predict the convergent behavior of the learned representations. Second, based on the actual representation objective, we draw a connection between METRA and contrastive learning. In Sec. 6, we conduct experiments to verify that METRA learns optimal representations in practice and that they bear resemblance to contrastive representations.
@@
where $f:{\mathcal{S}}\times{\mathcal{S}}\times{\mathcal{Z}}\mapsto\mathbb{R}$ is the critic function (Ma & Collins, 2018; Poole et al., 2019; Eysenbach et al., 2022; Zheng et al., 2024). Since the critic function $f$ takes arbitrary functional form, one can choose to parameterize $f$ as the inner product between the difference of transition representations and the latent skill, i.e. $f(s,s^{\prime},z)=(\phi(s^{\prime})-\phi(s))^{\top}z$. This yields a specific lower bound:
@@
The METRA representation objective is equivalent to a second-order Taylor approximation of $I^{\beta}(S,S^{\prime};Z)$, i.e., $L(\phi,\lambda_{0}(d))\approx\text{LB}^{\beta}(\phi)$.
@@ 4.2 Connecting METRA’s Actor Objective with an Information Bottleneck
This section discusses the actor objective used in METRA. We first clarify the distinction between the actor objective of METRA and those used in prior methods, helping to identify a term that discourages exploration. Removing this anti-exploration term results in covering a larger proposition of the state space while learning distinguishable skills. We then relate this anti-exploration term to estimating another mutual information, drawing a connection between the entire METRA actor objective and a variant of the information bottleneck (Tishby et al., 2000; Alemi et al., 2017).
@@
where $\text{LB}^{\pi}(\phi)$, $\text{LB}^{\pi}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+}}(\phi)$, and $\text{LB}^{\pi}_{-}(\phi)$ are under the target policy $\pi$ instead of the behavioral policy $\beta$. The SOTA performance of METRA and the divergence between the functional form of the actor objective (positive term) and the representation objective (positive and negative terms) suggests that $\text{LB}^{\pi}_{-}(\phi)$ may be a term discouraging exploration. Intuitively, removing this anti-exploration term boosts the learning of diverse skills. We will empirically study the effect of the anti-exploration term in Sec. 6.3 and provide theoretical interpretations next.
Our understanding of the anti-exploration term $\text{LB}^{\pi}_{-}(\phi)$ relates it to a resubstituion estimation of the differential entropy $h^{\pi}(\phi(S^{\prime})-\phi(S))$ in the representation space (see Appendix A.4 for details), i.e., $\text{LB}^{\pi}_{-}(\phi)=\hat{h}^{\pi}(\phi(S^{\prime})-\phi(S))$. Note that this entropy is different from the entropy of states $h^{\pi}(S)$, indicating that we want to minimize the entropy of difference of representations $\phi(s^{\prime})-\phi(s)$ to encourage exploration. There are two underlying reasons for this (seemly counterintuitive) purpose: METRA aims to (1) constrain the expected L2 distance of difference of representations $\phi(s^{\prime})-\phi(s)$ (Eq. 5) and (2) push difference of representations $\phi(s^{\prime})-\phi(s)$ towards skills $z$ sampled from $\textsc{Unif}({\mathbb{S}}^{d-1})$. Nonetheless, this relationship allows us to further rewrite the anti-exploration term $\text{LB}^{\pi}_{-}(\phi)$ as an estimation of the mutual information $I^{\pi}(\phi(S^{\prime})-\phi(S^{\prime});S,S^{\prime})$, connecting the METRA actor objective to an information bottleneck:
@@ 5.2 Learning a Policy with Successor Features
Algorithm Summary. In Alg. 1, we summarize CSF, our new algorithm. [^4] Code: https://github.com/Princeton-RL/contrastive-successor-features Starting from an existing MISL algorithm (e.g., DIAYN (Eysenbach et al., 2019) and METRA (Park et al., 2024)), implementing our algorithm requires making three simple changes: (1) learning state representations $\phi_{\theta}$ by minimizing an InfoNCE loss (excluding positive pairs in the denominator) between pairs of $(s,s^{\prime})$ and $z$(2) using a critic $\psi_{\omega}$ with $d$-dimensional outputs and replacing the scalar reward with the vector $\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)$(3) sampling the action $a$ from the policy $\phi_{\eta}$ to maximize the inner product $\psi_{\omega}(s,a,z)^{\top}z$.
@@ 6.2 METRA Learns Contrastive Representations
Figure 2: Histograms of METRA representations. (a) The expected distance of representations converges to $1.0$, helping to explain what objective METRA’s representations are optimizing. (b) Given a latent skill, the conditional difference in representations ($\phi(s^{\prime})-\phi(s)\mid z$) converges to an isotropic Gaussian distribution. (c) Taking the marginal over latent skills, the normalized difference in representations $\left(\frac{(\phi(s^{\prime})-\phi(s))}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}\right)$ converges to a $\textsc{Unif}({\mathbb{S}}^{d-1})$. These observations are consistent with our theoretical analysis (Cor. 1) suggesting that METRA is performing a form of contrastive learning.
@@
To answer this question, we reuse the trained algorithm in Sec. 6.1 and visualize two important statistics: (1) the conditional differences in representations $\phi(s^{\prime})-\phi(s)-z$ and (2) the normalized marginal differences in representations $(\phi(s^{\prime})-\phi(s))/\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}$. The resulting histograms (Fig. 2(b) &  2(c)) indicate that the conditional differences in representations $\phi(s^{\prime})-\phi(s)-z$ converges to an isotropic Gaussian in distribution while the normalized marginal differences in representations $(\phi(s^{\prime})-\phi(s))/\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}$ converges to a uniform distribution on the $d$-dimensional unit hypersphere ${\mathbb{S}}^{d-1}$ in distribution. Prior work (Wang & Isola, 2020) has shown that representations derived from contrastive learning preserves properties similar to these observations. We conjecture that maximizing the contrastive lower bound on $I^{\beta}(S,S^{\prime};Z)$ directly has the same effect as maximizing the METRA representation objective. See Appendix A.8 for formal claims and connections.

### v3

@@ 1 Introduction
Self-supervised learning has had a large impact on areas of machine learning ranging from audio processing (Oord et al., 2016; 2018) or computer vision (Radford et al., 2021; Chen et al., 2020) to natural language processing (Devlin et al., 2019; Radford & Narasimhan, 2018; Radford et al., 2019; Brown, 2020). In the reinforcement learning (RL) domain, the “right” recipe to apply self-supervised learning is not yet clear. Several self-supervised methods for RL directly apply off-the-shelf methods from other domains, such as masked autoencoding (Liu et al., 2022), but have achieved limited success so far. Other methods design self-supervised routines more specifically built for the RL setting (Burda et al., 2019; Pathak et al., 2017; Eysenbach et al., 2019; Sharma et al., 2020; Pong et al., 2020). We will focus on the skill learning methods, which aim to learn a set of diverse and distinguishable behaviors (skills) without an external reward function. This objective is typically formulated as maximizing the mutual information between skills and states (Gregor et al., 2016; Eysenbach et al., 2019), namely mutual information skill learning (MISL). However, some promising recent advances in skill learning methods build on other intuitions such as Lipschitz constraints (Park et al., 2022) or transition distances (Park et al., 2023). This paper focuses on determining whether the good performance of those recent methods can still be explained within the well-studied framework of mutual information maximization.
@@
Building upon our new interpretations of METRA, we propose a simpler and competitive MISL algorithm called Contrastive Successor Features (CSF). First, CSF learns state representations by directly optimizing a contrastive lower bound on mutual information, preventing the dual gradient descent procedure adopted by METRA. Second, while any off-the-shelf RL algorithm (e.g., SAC (Haarnoja et al., 2018)) is applicable, CSF instead learns a policy by leveraging successor features of linear rewards defined by the learned representations. Experiments on six continuous control tasks show that CSF is comparable with METRA, as evaluated on exploration performance and on downstream tasks. Furthermore, ablation studies suggest that rewards derived from the information bottleneck, as well as a specific parameterization of representations, are key to good performance.
@@
Figure 1: From METRA to MISL. (Left) METRA argues that optimizing a Wasserstein distance is superior to using mutual information. (Right) Through careful analysis, we show METRA still bears striking similarities to MISL algorithms, which allows us to develop a new MISL algorithm (CSF) that matches the performance of METRA while retaining the theoretical properties associated with MI maximization.
@@ 2 Related Work
Unsupervised skill discovery. Our work builds upon prior methods that perform unsupervised skill discovery. Prior work has achieved this aim by maximizing lower bounds (Tschannen et al., 2020; Poole et al., 2019) of different mutual information formulations, including diverse and distinguishable skill-conditioned trajectories (Li et al., 2023; Eysenbach et al., 2019; Hansen et al., 2020; Laskin et al., 2022; Strouse et al., 2022), intrinsic empowerment (Mohamed & Jimenez Rezende, 2015; Choi et al., 2021), distinguishable termination states (Gregor et al., 2016; Warde-Farley et al., 2019; Baumli et al., 2021), entropy bonus (Florensa et al., 2016; Lee et al., 2019; Shafiullah & Pinto, 2022), predictable transitions (Sharma et al., 2020; Campos et al., 2020), etc. Among those prior methods, perhaps the most related works are CIC (Laskin et al., 2022) and VISR (Hansen et al., 2020). We will discuss the difference between them and our method in Sec. 5. Another line of unsupervised skill learning methods uses ideas other than mutual information maximization, such as Lipschitz constraints (Park et al., 2022), MDP abstraction (Park & Levine, 2023), model learning (Park et al., 2023), and Wasserstein distance (Park et al., 2024; He et al., 2022). Our work will analyze the state-of-the-art method named METRA (Park et al., 2024) that builds on the Wasserstein dependency measure (Ozair et al., 2019), provide an alternative explanation under the well-studied MISL framework, and ultimately develop a simpler method.
@@
Contrastive learning. Contrastive learning has achieved great success for representation learning in natural language processing and computer vision (Radford et al., 2021; Chen et al., 2020; Gao et al., 2021; Sohn, 2016; Chopra et al., 2005; Oord et al., 2018; Gutmann & Hyvärinen, 2010; Ma & Collins, 2018; Tschannen et al., 2020). These methods aim to push together the representations of positive pairs drawn from the joint distribution, while pushing away the representations of negative pairs drawn from the marginals (Ma & Collins, 2018; Oord et al., 2018). In the domain of RL, contrastive learning has been used to define auxiliary representation learning objectives for control (Laskin et al., 2020; Yarats et al., 2021), solve goal-conditioned RL problems (Zheng et al., 2024; Eysenbach et al., 2022; 2020; Ma et al., 2023), and derive representations for skill discovery (Laskin et al., 2022). Prior work has also provided theoretical analysis for these methods from the perspective of mutual information maximization (Poole et al., 2019; Tschannen et al., 2020) and the geometry of learned representations (Wang & Isola, 2020). Our work will combine insights from both angles to analyze METRA and show its relationship to contrastive learning, resulting in a new skill learning method.
@@
Successor features. Our work builds on successor representations (Dayan, 1993), which encode the discounted state occupancy measure of policies. Prior work has shown these representations can be learned on high-dimensional tasks (Kulkarni et al., 2016; Zhang et al., 2017) and help with transfer learning (Barreto et al., 2017). When combined with universal value function approximators (Schaul et al., 2015), these representations generalize to universal successor features, which estimate a value function for any reward under any policy (Borsa et al., 2019). While prior methods have combined successor feature learning with mutual information skill discovery for fast task inference (Hansen et al., 2020; Liu & Abbeel, 2021), we instead use successor features to replace $Q$ estimation after learning state representations (Sec. 5).
@@ 3 Preliminaries
where $p^{\pi}(s_{+}=s\mid z)$ is the discounted state occupancy measure (Ho & Ermon, 2016; Nachum et al., 2019; Eysenbach et al., 2022; Zheng et al., 2024) of policy $\pi$ conditioned on skill $z$, and $p^{\pi}(s^{\prime}\mid s,z)$ is the state transition probability given policy $\pi$ and skill $z$. This optimization problem can be cast as an iterative min-max optimization problem by first choosing a variational distribution $q(z\mid s,s^{\prime})$ to fit the historical posterior $p^{\beta}(z\mid s,s^{\prime})$, which is an approximation of $p^{\pi}(z\mid s,s^{\prime})$, and then choosing policy $\pi$ to maximize discounted return defined by the intrinsic reward $\log q(z\mid s,s^{\prime})$:
@@
where $k$ indicates the number of updates. See Appendix A.1 for a detailed discussion.
@@
METRA. Maximizing the mutual information between states and latent skills $I(S;Z)$ only encourages an agent to find discriminable skills, while the algorithm might fail to prioritize state space coverage (Park et al., 2024; 2022). A prior state-of-the-art method, METRA (Park et al., 2024), proposes to solve this problem by learning representations of states $\phi:{\mathcal{S}}\mapsto\mathbb{R}^{d}$ via maximizing the Wasserstein dependency measure (WDM) (Ozair et al., 2019) between states and skills $I_{\mathcal{W}}(S;Z)$. Specifically, METRA chooses to enforce the 1-Lipschitz continuity of $\phi$ under the temporal distance metric, resulting in a constrained optimization problem for $\phi$:
@@ 4 Understanding the Prior Method
In this section, we reinterpret METRA through the lens of MISL, showing that:
@@
    The METRA actor objective is equivalent to a mutual information lower bound plus an extra term. This extra term is related to an information bottleneck (Tishby et al., 2000; Alemi et al., 2017), and our experiments will show that it is important for exploration. See Sec. 4.2.
@@ 4.1 Connecting METRA’s Representation Objective and Contrastive Learning
Our understanding of METRA starts by interpreting the representation objective of METRA as a contrastive loss. This interpretation proceeds in two steps. First, we focus on understanding the actual representation objective of METRA, aiming to predict the convergent behavior of the learned representations. Second, based on the actual representation objective, we draw a connection between METRA and contrastive learning. In Sec. 6, we conduct experiments to verify that METRA learns optimal representations in practice and that they bear resemblance to contrastive representations.
@@
where $f:{\mathcal{S}}\times{\mathcal{S}}\times{\mathcal{Z}}\mapsto\mathbb{R}$ is the critic function (Ma & Collins, 2018; Poole et al., 2019; Eysenbach et al., 2022; Zheng et al., 2024). Since the critic function $f$ takes an arbitrary functional form, one can choose to parameterize $f$ as the inner product between the difference of transition representations and the latent skill, i.e., $f(s,s^{\prime},z)=(\phi(s^{\prime})-\phi(s))^{\top}z$. This yields a specific lower bound:
@@
The METRA representation objective is equivalent to a second-order Taylor approximation of the contrastive lower bound on $I^{\beta}(S,S^{\prime};Z)$, i.e., $L(\phi,\lambda_{0}(d))\approx\text{LB}^{\beta}(\phi)$.
@@ 4.2 Connecting METRA’s Actor Objective with an Information Bottleneck
This section discusses the actor objective used in METRA. We first clarify the distinction between the actor objective of METRA and those used in prior methods, helping to identify a term that discourages exploration. Removing this anti-exploration term results in covering a larger proportion of the state space while learning distinguishable skills. We then relate this anti-exploration term to estimating another mutual information, drawing a connection between the entire METRA actor objective and a variant of the information bottleneck (Tishby et al., 2000; Alemi et al., 2017).
@@
where $\text{LB}^{\pi}(\phi)$, $\text{LB}^{\pi}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+}}(\phi)$, and $\text{LB}^{\pi}_{-}(\phi)$ are under the target policy $\pi$ instead of the behavioral policy $\beta$. The SOTA performance of METRA and the divergence between the functional form of the actor objective (positive term) and the representation objective (positive and negative terms) suggest that $\text{LB}^{\pi}_{-}(\phi)$ may be a term discouraging exploration. Intuitively, removing this anti-exploration term boosts the learning of diverse skills. We will empirically study the effect of the anti-exploration term in Sec. 6.3 and provide theoretical interpretations next.
Our understanding of the anti-exploration term $\text{LB}^{\pi}_{-}(\phi)$ relates it to a resubstituion estimation of the differential entropy $h^{\pi}(\phi(S^{\prime})-\phi(S))$ in the representation space (see Appendix A.4 for details), i.e., $\text{LB}^{\pi}_{-}(\phi)=\hat{h}^{\pi}(\phi(S^{\prime})-\phi(S))$. Note that this entropy is different from the entropy of states $h^{\pi}(S)$, indicating that we want to minimize the entropy of the difference of representations $\phi(s^{\prime})-\phi(s)$ to encourage exploration. There are two underlying reasons for this (seemly counterintuitive) purpose: METRA aims to (1) constrain the expected L2 distance of difference of representations $\phi(s^{\prime})-\phi(s)$ (Eq. 5) and (2) push difference of representations $\phi(s^{\prime})-\phi(s)$ towards skills $z$ sampled from $\textsc{Unif}({\mathbb{S}}^{d-1})$. Nonetheless, this relationship allows us to further rewrite the anti-exploration term $\text{LB}^{\pi}_{-}(\phi)$ as an estimation of the mutual information $I^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S^{\prime}))$, connecting the METRA actor objective to an information bottleneck:
@@ 5.2 Learning a Policy with Successor Features
Algorithm Summary. In Alg. 1, we summarize CSF, our new algorithm. [^4] Code: https://github.com/Princeton-RL/contrastive-successor-features Starting from an existing MISL algorithm (e.g., DIAYN (Eysenbach et al., 2019) and METRA (Park et al., 2024)), implementing our algorithm requires making three simple changes: (1) learning state representations $\phi_{\theta}$ by minimizing an InfoNCE loss (excluding positive pairs in the denominator) between pairs of $(s,s^{\prime})$ and zz(2) using a critic $\psi_{\omega}$ with $d$-dimensional outputs and replacing the scalar reward with the vector $\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)$(3) sampling the action $a$ from the policy $\pi_{\eta}$ to maximize the inner product $\psi_{\omega}(s,a,z)^{\top}z$, where $\theta$, $\omega$, and $\eta$ denote parameters of neural networks.
@@ 6.2 METRA Learns Contrastive Representations
Figure 2: Histograms of METRA representations. (a) The expected distance of representations converges to $1.0$, helping to explain what objective METRA’s representations are optimizing. (b) Given a latent skill, the conditional difference in representations ($\phi(s^{\prime})-\phi(s)\mid z$) converges to an isotropic Gaussian distribution. (c) Taking the marginal over latent skills, the normalized difference in representations $\left(\frac{(\phi(s^{\prime})-\phi(s))}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}\right)$ converges to a $\textsc{Unif}({\mathbb{S}}^{d-1})$. These observations are consistent with our theoretical analysis (Cor. 1), suggesting that METRA is performing a form of contrastive learning.
@@
To answer this question, we reuse the trained algorithm in Sec. 6.1 and visualize two important statistics: (1) the conditional differences in representations $\phi(s^{\prime})-\phi(s)-z$ and (2) the normalized marginal differences in representations $(\phi(s^{\prime})-\phi(s))/\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}$. The resulting histograms (Fig. 2(b) &  2(c)) indicate that the conditional differences in representations $\phi(s^{\prime})-\phi(s)-z$ converges to an isotropic Gaussian in distribution while the normalized marginal differences in representations $(\phi(s^{\prime})-\phi(s))/\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}$ converges to a uniform distribution on the $d$-dimensional unit hypersphere ${\mathbb{S}}^{d-1}$ in distribution. Prior work (Wang & Isola, 2020) has shown that representations derived from contrastive learning preserve properties similar to these observations. We conjecture that maximizing the contrastive lower bound on $I^{\beta}(S,S^{\prime};Z)$ directly has the same effect as maximizing the METRA representation objective. See Appendix A.8 for formal claims and connections.